Cryptology Problems

To be discussed Wednesday, April 15, 1998.

1. Do problems 6.1.a, 6.1.b, and all of 6.5 in the textbook.

2. (From Salomaa’s textbook) In the ElGamal Signature Scheme, what is the effect on ecryption, decryption, and security if α is not primitive element? (Note that α should be a primitive element, so this would change the scheme somewhat.)